# Existence of Minimal Surfaces in Infinite Volume Hyperbolic $3$-manifolds

### v1

@@ 7.1. Summary of Existence Results
  • $\mathcal{M}$ is geometrically finite: If $\mathcal{M}$ is topologically not a handlebody, then there exists a finite area, complete, embedded, minimal surface in $\mathcal{M}$ (6.11 and 6.14 when $\mathcal{M}$ is quasi-Fuchsian ).
  • $\mathcal{M}$ is geometrically infinite: If $\mathcal{M}$ is not a handlebody nor a product, and contains a convex barrier surface in one of the ends, then there exists a finite area, complete, embedded, minimal surface in $\mathcal{M}$ (6.10).
@@ 7.2. Open Questions
In this part, we discuss some open questions and potential approaches. For doubly degenerate manifolds with bounded geometry, we proved a dichotomy that either any such manifold contains a closed minimal surface, or there must be a doubly degenerate manifold which contains a minimal foliation. Notice that this result is closely related with another important question posed by Thurston decades ago, also known as Anderson’s Conjecture (for a negative answer) [And83]:
@@
This is a decades-old famous question in the domain which relates to our dichotomy result through Virtual Fibering Theorem due to Agol [Ago13]. With this result, any closed hyperbolic $3$-manifold with minimal foliation has a finite cover which fibers over a circle. Hence, the infinite cover of such a fibering hyperbolic manifold will give a doubly degenerate manifold with bounded geometry as stated in our dichotomy (Theorem 4.3). A negative answer to 7.1 does not rule out the second case in the dichotomy, as there might be foliations in doubly degenerate manifolds, which do not descend to closed manifolds.
@@
Our proof of 4.3 used the assumption of bounded geometry in an essential way. If one wants to adapt the arguments to apply to the unbounded geometry case, allowing arbitrarily short geodesics, a more complicated analysis is needed to find a minimizer for the energy functional. This involves serious difficulties first in the harmonic maps part of the argument, where one would need to consider harmonic maps from possibly cusped surfaces into targets without bounded geometry. There are also complications in applying the theory of convergence for Kleinian groups, which is much more involved if one drops the bounded geometry assumption. Even so, we believe that one could obtain similar results in this setting, and that it is an interesting topic for future work.
The next question is about zero-infinity dichotomy for non-cusped hyperbolic $3$-manifolds.
@@
Let $\mathcal{M}$ be a geometrically infinite hyperbolic $3$-manifold. Then, is it true that if $\mathcal{M}$ contains infinitely many finite area, complete, embedded, minimal surfaces?
@@
Identify all Schottky manifolds without a closed embedded minimal surface.
We also ask whether Theorem 4.5 and Corollary 4.6 on upper bounds for the number of disjoint stable minimal surfaces in the convex core of a quasi-Fuchsian manifold can be generalized.
@@
For manifolds with rank-1 cusps, we pose two questions. The first one is about constructing min-max surfaces in these manifolds.
@@
In 6.10, if all the barrier surfaces are mean-concave, is it possible to establish a min-max surface in $\mathcal{M}$?
This question is closely connected to adapting Song’s min-max result [Son23a, Corollary 2.2], which establishes the existence of a min-max surface when $\mathcal{M}$ contains a compact mean-concave domain. Our construction in Section 6.1 yields a finite-volume mean-concave domain. It may be possible to refine this construction to obtain a compact mean-concave domain, thereby enabling the application of [Son23a, Corollary 2.2] to produce a min-max surface with finite area.
The second question regarding manifolds with rank-1 cusps is another open case, analogous to the earlier product case:
Question 7.7.
Let $\mathcal{M}$ be a geometrically infinite hyperbolic $3$-manifold with =×$\mathcal{M}=\mathcal{F}\times\mathbf{R}$ where $\mathcal{F}$ is a punctured surface. Does $\mathcal{M}$ contains a finite area, complete, embedded, minimal surface?

### v2

@@ 7.1. Summary of Existence Results
  • $\mathcal{M}$ contains a geometrically finite geometric end: If $\mathcal{M}$ is topologically not a handlebody nor product, then there exists a finite area, complete, embedded, minimal surface in $\mathcal{M}$ (6.2, and when $\mathcal{M}$ is quasi-Fuchsian , see [HLS23]).
@@ 7.2. Open Questions
In this part, we discuss some open questions and potential approaches. For doubly degenerate manifolds with bounded geometry, we proved that either any such manifold contains a closed minimal surface, or if not there must be a doubly degenerate manifold which has a minimal foliation. Notice that this result is closely related with another important question posed by Thurston decades ago, also known as Anderson’s Conjecture (for a negative answer) [And83]:
@@
This is a decades-old famous question in the domain which relates to our alternative result through Virtual Fibering Theorem due to Agol [Ago13]. With this result, any closed hyperbolic $3$-manifold with minimal foliation has a finite cover which fibers over a circle. Hence, the infinite cover of such a fibering hyperbolic manifold will give a doubly degenerate manifold with bounded geometry as in the statement of Theorem 4.3. A negative answer to 7.1 does not rule out the second case in the alternative, as there might be foliations in doubly degenerate manifolds, which do not descend to closed manifolds.
@@
Our proof of 4.3 used the bounded geometry assumption in an essential way. If one wants to adapt the arguments to apply to the unbounded geometry case, allowing arbitrarily short geodesics, a more complicated analysis is needed to find a minimizer for the energy functional. This involves serious difficulties first in the harmonic maps part of the argument, where one would need to consider harmonic maps from possibly cusped surfaces into targets without bounded geometry. There are also complications in applying the theory of convergence for Kleinian groups, which is much more involved if one drops the bounded geometry assumption. Even so, we believe that one could obtain similar results in this setting, and that it is an interesting topic for future work.
The next question concerns the existence problem for minimal surfaces in general hyperbolic 3-manifolds. All the known examples of hyperbolic $3$-manifolds with no finite area, embedded minimal surfaces are geometrically finite (5.3). Furthermore, our dichotomy (3.5) shows that the existence of one saddle minimal surface implies the existence of infinitely many finite area, embedded minimal surfaces. Hence, we pose the following question.
@@
Let $\mathcal{M}$ be a geometrically infinite hyperbolic $3$-manifold with no cusps. Then, is it true that $\mathcal{M}$ contains infinitely many finite area, complete, embedded, minimal surfaces?
@@
Characterize the Schottky manifolds that have no closed embedded minimal surface.
We also ask whether Theorem 4.5 and Corollary 4.6 on upper bounds for the number of disjoint stable minimal surfaces in the convex core of a quasi-Fuchsian manifold can be strengthened.
@@
In the setting of hyperbolic 3–manifolds with rank-1 cusps, although we addressed an important special case, the general problem remains wide open.
@@
Let $\mathcal{M}$ be a hyperbolic $3$-manifold with rank-1 cusps. Does $\mathcal{M}$ contain a finite area, complete, embedded minimal surface?
A promising strategy is to generalize the shrinkwrapping techniques from [Cos21] to this cusp setting. However, the absence of a uniform positive lower bound on the injectivity radius makes several steps of that construction highly delicate.